$SPAD/src/input atansqrt.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)set break resume
)spool atansqrt.output
)set message test on
)set message auto off
)clear all

———-

given — * —

--S 1 of 3
z:=atan sqrt ((1-cos x)/(1+cos x))
--R
--R
--R              +------------+
--R              |- cos(x) + 1
--R   (1)  atan( |------------ )
--R             \| cos(x) + 1
--R                                                    Type: Expression(Integer)
--E 1

———-

this should generate x/2 — * —

--S 2 of 3
integrate(differentiate(z,x),x)
--R
--R
--R        x
--R   (2)  -
--R        2
--R                                         Type: Union(Expression(Integer),...)
--E 2

———-

so should this be x/2 — * —

--S 3 of 3
rootSimp(normalize(z))
--R
--R
--R        x
--R   (3)  -
--R        2
--R                                                    Type: Expression(Integer)
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

5